### 31 Jan 2024 v1

# On the optimal objective value of random linear programs

Abstract
We consider the problem of maximizing $\langle c,x\rangle$ subject to the constraints $Ax\leq\mathbf{1}$, where $x\in\mathbb{R}^{n}$, $A$ is an $m\times n$ matrix with mutually independent symmetric subgaussian entries of unit variance, and $c$ is a normalized cost vector. In the regime $n\to\infty$, $\frac{m}{n}\to\infty$, and under the assumption that $c$ is incompressible, we obtain sharp lower bounds on the optimal objective value $z^{*}$ of the linear program. In the context of high-dimensional convex geometry, our findings imply that the spherical mean width ${\mathcal{W}}(P)$ of the random convex polyhedron $P=\{x\in\mathbb{R}^{n}:\;Ax\leq\mathbf{1}\}$ satisfies
$\liminf\limits_{n\to\infty}\sqrt{2\log(m/n)}\,{\mathcal{W}}(P)\geq 2$
almost surely, with the matching upper bound available in the Gaussian setting.
We provide extensive numerical experiments as supporting data for the theoretical predictions. Further, we carry out numerical studies for limiting distributions and standard deviations of $z^{*}$ and ${\mathcal{W}}(P)$, and formulate several conjectures regarding their asymptotics.
1 Introduction
1.1 Problem setup and motivation
Consider the linear program (LP)
$\begin{split}&\text{maximize}\ \ \langle c,x\rangle\\ &\text{subject to }\ Ax\leq\mathbf{1}\\ &x\in\mathbb{R}^{n}\end{split}$ (1)
where the entries of the coefficient matrix $A$ are mutually independent random variables, and ${\bf 1}$ denotes the vector of all ones. Note that $x$ is a vector of free variables. In this note we deal with two settings: either $c$ is a non-random unit vector or $c$ is random uniformly distributed on the unit Euclidean sphere. For any realization of $A$ the above LP is feasible since the zero vector satisfies all the constraints. We denote the value of the objective function of (1) by $z=z(x)=\langle c,x\rangle$, the optimal objective value by $z^{*}$, and any optimal solution by $x^{*}$.
The main problem we would like to address here is the following: Under the assumption that the number of constraints is significantly larger than the number of variables, what is the magnitude of the optimal objective value?
We provide theoretical guarantees on the optimal objective value $z^{*}$ in the asymptotic regime $n\to\infty$, $\frac{m}{n}\to\infty$, and carry out empirical studies which confirm that medium size random LPs (with the number of constraints in the order of a few hundreds to a few thousands) closely follow theoretical predictions. The numerical experiments further allow us to formulate a few conjectures related to the distribution of the optimal objective value, which we hope to address in future works.
Our interest in LP (1) is twofold.
  • (i)
    The LP (1) with i.i.d random coefficients can be viewed as the simplest model of an “average-case” linear program. LP of this form has been considered in the mathematical literature in the context of estimating the average-case complexity of the simplex method [8, 9, 10] (see Subsection 1.4 below for a more comprehensive literature review). Despite significant progress in that direction, fundamental questions regarding properties of random LPs remain open. The results obtained in this note contribute to better understanding of the geometry of random feasible regions defined as intersections of independent random half-spaces. We view our work as a step towards investigating random LPs allowing for sparse, inhomogeneous and correlated constraints, which would serve as more realistic models of linear programs occurring in practice.
  • (ii)
    Assuming that the cost vector $c$ is uniformly distributed on the unit sphere and is independent from $A$, the quantity
    $2\,{\mathbb{E}}_{c}\,z^{*}=2\,{\mathbb{E}}_{c}\,\max\big{\{}\langle c,x\rangle:\;Ax\leq{\bf 1}\big{\}}$
    (where ${\mathbb{E}}_{c}$ is the expectation taken with respect to the randomness of $c$) is the spherical mean width ${\mathcal{W}}(P)$ of the random polyhedron $P=\{x\in\mathbb{R}^{n}:\;Ax\leq\mathbf{1}\}$. The mean width is a classical parameter associated with a convex set; in particular, ${\mathcal{W}}(P)=M(P^{\circ})$, where $P^{\circ}$ is the polar body for $P$ defined as the convex hull of the rows of $A$, and $M(\cdot)$ is the average value of the Minkowski functional for $P^{\circ}$ on the unit Euclidean sphere,
    $M(P^{\circ})={\mathbb{E}}_{c}\,\|c\|_{P^{\circ}}={\mathbb{E}}_{c}\,\inf\big{\{}\lambda>0:\,\lambda^{-1}c\in P^{\circ}\big{\}}.$
    The mean width plays a key role in the study of projections of convex bodies and volume estimates (see, in particular, [50, Chapters 9–11] and [40]). The existing body of work (which we are going to revise in Subsection 1.4) has focused on the problem of estimating the mean width up to a constant multiple. To the best of our knowledge, our paper is the first to provide sharp asymptotics of ${\mathcal{W}}(P)$ in the regime of parameters mentioned above.
1.2 Notation
In the course of the note we use the letter $K$ (with subscripts) for constant numbers and $O(\cdot)$, $o(\cdot)$, $\Omega(\cdot)$, $\omega(\cdot)$ for standard asymptotic notations. Specifically, for positive functions $f$ and $g$ defined on positive real numbers, $f(x)=O\big{(}g(x)\big{)}$ if there exists $K>0$ and a real number $x_{0}$ such that $f(x)\leq Kg(x)$ for all $x\geq x_{0}$, and $f(x)=o\big{(}g(x)\big{)}$ if $\lim_{x\rightarrow\infty}\frac{f(x)}{g(x)}=0$. Further,
$f(x)=\Omega\big{(}g(x)\big{)}\ \ \text{if and only if}\ \ g(x)=O\big{(}f(x)\big{)},$
$f(x)=\omega\big{(}g(x)\big{)}\ \ \text{if and only if}\ \ g(x)=o\big{(}f(x)\big{)}.$
Definition 1 (Sparse vectors).
A vector $v$ in $\mathbb{R}^{n}$ is $s$–sparse for some $s>0$ if the number of non-zero components in $v$ is at most $s$.
Definition 2 (Compressible and incompressible vectors, [42]).
Let $v$ be a unit vector in $\mathbb{R}^{n}$, and let $\delta,\rho\in(0,1)$. The vector $v$ is called $(\delta,\rho)$–compressible if the Euclidean distance from $v$ to the set of $\delta n$–sparse vectors is at most $\rho$, that is, if there is a $\delta n$–sparse vector $z$ such that $\|v-z\|_{2}\leq\rho$. Otherwise, $v$ is $(\delta,\rho)$–incompressible.
Observe that a vector $v$ is $(\delta,\rho)$–compressible if and only if the sum of squares of its $\lfloor\delta n\rfloor$ largest (by the absolute value) components is at least $1-\rho^{2}$.
Definition 3.
A random variable $\xi$ is $K$–subgaussian if
${\mathbb{E}}\,\exp(\xi^{2}/K^{2})\leq 2,\ \text{for some }\ K>0.$
The smallest $K$ satisfying the above inequality is the subgaussian moment of $\xi$ (see, for example, [50, Chapter 2]). We will further call a random vector with mutually independent components $K$–subgaussian if every component is $K$–subgaussian. Examples of subgaussian variables include Rademacher (i.e symmetric ±$\pm 1$) random variables, and, more generally, any bounded random variables. The Rademacher distribution is of interest to us since a matrix with i.i.d ±$\pm 1$ entries produces a simplest model of a normalized system of constraints with discrete bounded coefficients.
1.3 Theoretical guarantees
In this paper, we study the asymptotics of the optimal objective value of (1) in the regime when the number of free variables $n$ tends to infinity. We will further assume that the number of constraints $m=m(n)$ is infinitely large compared to $n$, i.e $\lim\limits_{n\to\infty}\frac{m}{n}=\infty$ (while the setting $\frac{m}{n}=O(1)$ is of significant interest, it is not covered by our argument).
Theorem 4 (Lower bound on $z^{*}$).
Let $K>0$. Assume that $\lim\limits_{n\to\infty}\frac{m}{n}=\infty$ and that for each $n$, the entries of the $m\times n$ coefficient matrix $A=A(n)$ are mutually independent symmetric $K$–subgaussian variables of unit variance. Assume further that the non-random unit cost vectors $c=c(n)$ satisfy
for every $\kappa>0$ there is $n_{0}(\kappa)\in\mathbb{N}$ such that for all $n\geq n_{0}(\kappa)$,
$c=c(n)$ is $(\kappa^{-1}\log(m/n)/n,1-\kappa)$–incompressible.
Let $z^{*}=z^{*}(n)$ be the optimal objective value of (1). Then for any constant $\varepsilon,K^{\prime}>0$ and all sufficiently large $n$, we have
${\mathbb{P}}\big{\{}\sqrt{2\log(m/n)}\,z^{*}\geq 1-\varepsilon\big{\}}\geq 1-n^{-K^{\prime}}.$
In particular, $\liminf\limits_{n\to\infty}\sqrt{2\log(m/n)}\,z^{*}\geq 1$ almost surely.
Observe that the lower bound in Theorem 4 is not influenced by a specific choice of the subgaussian distribution of the matrix entries, and is only determined by the aspect ratio $m/n$. We refer to Section 5.2 for a discussion of the above assumptions on the cost vectors and related numerical experiments. Although the assumptions on the cost vectors in Theorem 4 may appear excessively technical, we conjecture that they are optimal in the following sense: whenever there is $\kappa>0$ such that $c(n)$ is $(\kappa^{-1}\log(m/n)/n,1-\kappa)$–compressible for infinitely many $n\in\mathbb{N}$, there is a symmetric subgaussian variable $\xi$ of unit variance such that, assuming $A(n)$ has i.i.d entries equidistributed with $\xi$, we have $\liminf\limits_{n\to\infty}\sqrt{2\log(m/n)}\,z^{*}<1$ almost surely.
As a simple sufficient condition on the cost vectors which implies the assumptions in Theorem 4, we consider a bound on the $\|\cdot\|_{\infty}$–norms:
Corollary 5 (Lower bound on $z^{*}$ under sup-norm delocalization of the cost vectors).
Let $m=m(n)$ satisfy $\lim\limits_{n\to\infty}\frac{m}{n}=\infty$, and let matrices $A=A(n)$ be as in Theorem 4. Consider a sequence of non-random cost vectors $c=c(n)$ satisfying $\lim_{n\rightarrow\infty}\sqrt{\log(m/n)}\,\|c\|_{\infty}=0$. Then for any constant $\varepsilon,K^{\prime}>0$ and all sufficiently large $n$,
${\mathbb{P}}\big{\{}\sqrt{2\log(m/n)}\,z^{*}\geq 1-\varepsilon\big{\}}\geq 1-n^{-K^{\prime}},$
implying that $\liminf\limits_{n\to\infty}\sqrt{2\log(m/n)}\,z^{*}\geq 1$ almost surely.
As we mentioned before, the setting when the cost vector $c$ is uniform on the Euclidean sphere is of special interest in the field of convex geometry since the quantity ${\mathcal{W}}=2\,{\mathbb{E}}_{c}\,z^{*}$ (where ${\mathbb{E}}_{c}$ denotes the expectation taken with respect to $c$) is the spherical mean width of the random polyhedron $P=P(n)=\{x\in\mathbb{R}^{n}:\;Ax\leq\mathbf{1}\}$. As another corollary of Theorem 4, we obtain
Corollary 6 (The mean width: a sharp lower bound).
Let $m=m(n)$ and $A=A(n)$ be as in Theorem 4, and assume additionally that $m=\exp(o(n))$. Then the spherical mean width ${\mathcal{W}}(P)$ of the polyhedron $P=P(n)=\{x\in\mathbb{R}^{n}:\;Ax\leq\mathbf{1}\}$ satisfies
$\liminf\limits_{n\to\infty}\sqrt{2\log(m/n)}\,{\mathcal{W}}(P)\geq 2\quad\mbox{almost surely.}$.
As the next proposition shows the bound of Theorem 4 is best possible among distribution-independent estimates. We recall that the outer radius of a bounded convex set $S$ is the smallest $R>0$ such that $S$ is entirely contained in the Euclidean ball of radius $R$ centered at the origin. It is easy to see that the optimal objective value $z^{*}$ is bounded above by the outer radius of the feasible region, regardless of the choice of the unit cost vector $c$.
Proposition 7 (Outer radius of the feasible region in the Gaussian setting).
Let $m=m(n)$ be as in Theorem 4, and for every $n$ let $A$ be an $m\times n$ random matrix with i.i.d standard Gaussian entries. Then, denoting by $R$ the outer radius of the feasible region, for any constant $\varepsilon,K>0$ and all sufficiently large $n$, we have
${\mathbb{P}}\big{\{}\sqrt{2\log(m/n)}\,R\leq 1+\varepsilon\big{\}}\geq 1-n^{-K}.$
In particular, $\limsup\limits_{n\to\infty}\sqrt{2\log(m/n)}\,R\leq 1$ almost surely.
Note that together with Corollary 6, Proposition 7, implies
Corollary 8.
Let $m=m(n)$ and $A=A(n)$ be as in Theorem 4, and assume additionally that $m=\exp(o(n))$ and that the entries of $A$ are i.i.d standard Gaussians. Then the spherical mean width ${\mathcal{W}}(P)$ of the polyhedron $P=P(n)=\{x\in\mathbb{R}^{n}:\;Ax\leq\mathbf{1}\}$ satisfies
$\lim\limits_{n\to\infty}\sqrt{2\log(m/n)}\,{\mathcal{W}}(P)=2\quad\mbox{almostsurely.}$.
We plan to investigate non-Gaussian setting in future works.
1.4 Review of related literature
1.4.1 Convex feasibility problems
In the setting of Theorem 4, existence of a feasible vector $x$ with $\sqrt{2\log(m/n)}\,\langle c,x\rangle\geq 1-\varepsilon$ is equivalent to the statement that the intersection of random convex sets
$C_{i}:=\bigg{\{}y\in c^{\perp}:\;\langle{\rm row}_{i}(A),y\rangle\leq 1-\frac{1-\varepsilon}{\sqrt{2\log(m/n)}}\langle{\rm row}_{i}(A),c\rangle\bigg{\}},\quad 1\leq i\leq m,$
is non-empty. The sets can be naturally viewed as independent random affine subspaces of the hyperplane $c^{\perp}$ and, with that interpretation, the above question is a relative of the spherical perceptron problem, in which the object of interest is the intersection of random affine halfspaces of the form $\{y:\,\langle R_{i},y\rangle\leq-\kappa\}$, for independent random vectors $R_{1},\dots,R_{m}$ and a constant parameter $\kappa\in\mathbb{R}$ (see [46, 47] for a comprehensive discussion).
Our proof of Theorem 4 is based on constructing a feasible solution $x$ to the LP (1) satisfying $\sqrt{2\log(m/n)}\,\langle c,x\rangle\geq 1-\varepsilon$ via an iterative process. We start by defining a vector $x_{0}$ as a multiple of the cost vector $c$, with $\langle c,x_{0}\rangle\approx(2\log(m/n))^{-1/2}$. Our goal is to add a perturbation to $x_{0}$ which would restore the feasibility. We shall select a vector $x_{1}\in c^{\perp}$ which would repair the constraints violated by $x_{0}$, more specifically, such that $\langle{\rm row}_{i}(A),x_{0}+x_{1}\rangle\leq 1$ for all $i\in[m]$ with $\langle{\rm row}_{i}(A),x_{0}\rangle>1$. The vector $x_{1}$ may itself introduce some constraints’ violation; to repair those constraints, we will find another vector $x_{2}\in c^{\perp}$, and consider $x_{0}+x_{1}+x_{2}$, etc. At $k$–th iteration, the vector $x_{k}$ will be chosen as a linear combination of the matrix rows having large scalar products with $x_{k-1}$. The process continues for some finite number of steps $r$ which depends on the values of $m$ and $n$. We shall prove that the resulting vector $x:=x_{0}+x_{1}+\dots+x_{r}$ is feasible with high probability. Since the choice of $x_{1},\dots,x_{r}$ is restricted to the orthogonal complement $c^{\perp}$, the value of the objective at the vector $x_{0}+x_{1}+\dots+x_{r}$ is the same as for the vector $x_{0}$.
From the algorithmic perspective, the above is an example of iterative approach to the convex feasibility problems which have been actively studied in optimization literature. The problem is to construct or approximate a vector in the non-empty intersection of a given collection of convex sets. The special setting where the convex sets are affine hyperplanes or halfspaces, corresponding to solving systems of linear equations and inequalities, was first addressed by Kaczmarz [28], Agmon [1], Motzkin and Schoenberg [37]. We refer, among many others, to papers [48, 2, 45, 17, 38, 12, 34, 25, 31, 27, 15, 5, 35, 39, 24, 33, 51, 36, 52, 53] for more recent developments and further references. In particular, the work [2] introduced a family of block-iterative projection methods for solving the feasibility problems, when each update is computed as a function of projections onto a subset of constraints.
The update rule in the proof of Theorem 4 which we informally described above, falls under the umbrella of the block-iterative projection methods. At the same time, we emphasize that the goal of our construction is to verify that the intersection is non-empty with high probability rather than approximating a solution to a problem which is known to be feasible a priori. Since we aim to get asymptotically sharp lower bounds on $z^{*}$, we have to rely on high-precision analysis of constraint’s violations, not allowing for any significant losses in our estimates.
1.4.2 Random linear programs
Linear programs with either the coefficient matrix or the coefficient vector generated randomly, have been actively studied in various contexts:
  • Average-case analysis. As we mentioned above, the setting of random LPs with the Gaussian coefficient matrix (or, more generally, a matrix with independent rows having spherically symmetric distributions) was considered by Borgwardt [8, 9, 10] in the context of estimating the average-case complexity of the shadow simplex method. The average-case complexity of the Dantzig algorithm was addressed by Smale in [43] (see also [6]).
  • Smoothed analysis. The smoothed analysis of the simplex algorithm, when the coefficient matrix is a sum of an arbitrary deterministic matrix and a small Gaussian perturbation, was first carried out by Spielman and Teng in [44], with improved estimates obtained in later works [49, 13, 26].
  • Distribution of the optimal objective value under small random perturbations. The setting when the coefficient matrix, the RHS or the cost vector is subject to small random perturbations was considered, in particular, in [4, 41, 29, 32].
  • Random integer feasibility problems. We refer to [11] for the problem of integer feasibility in the randomized setting.
  • The optimal objective value of LPs with random cost vectors [16].
1.4.3 The mean width of random polyhedra
As we have mentioned at the beginning of the introduction, the mean width of random polytopes has been studied in the convex geometry literature, although much of the work deals with the mean width of the convex hulls of random vectors i.e of the polars to the random polyhedra considered in the present work. We refer, in particular, to [14, 3, 21] for related results.
In the context of intersections of random half-spaces, it was shown in [22, 23] that the outer radius of the symmetric random polyhedron $P=\{x\in\mathbb{R}^{n}:\;Ax\leq{\bf 1}\;\mbox{and}\;-Ax\leq{\bf 1}\}$, where the matrix $A$ is standard Gaussian, is bounded above by $O\big{(}\frac{1}{\sqrt{\log(m/n)}}\big{)}$ with high probability; as we noted before, this estimate easily translates into the corresponding upper bound for the mean width of $P$.
A two-sided bound on the mean width in the i.i.d non-Gaussian setting was derived in [30] (see also [20] for earlier results); specifically, it was shown that under the assumption that the entries of $A$ are i.i.d subgaussian variables of zero mean and unit variance, the mean width of the symmetric polyhedron $P=\{x\in\mathbb{R}^{n}:\;Ax\leq{\bf 1}\;\mbox{and}\;-Ax\leq{\bf 1}\}$ satisfies
$\frac{K_{1}}{\sqrt{\log(m/n)}}\leq{\mathcal{W}}(P)\leq\frac{K_{2}}{\sqrt{\log(m/n)}}$ (2)
with high probability, where the constants $K_{1},K_{2}$ depend on the subgaussian moment of the matrix entries. The argument in [30] heavily relies on results which introduce suboptimal constant multiples to the estimate of ${\mathcal{W}}(P)$, and we believe that it cannot be adapted to derive an asymptotically sharp version of (2) which is the main goal of the present work.
1.5 Organization of the paper
We derive the proof of Theorem 4 in Section 2. Corollaries 5 and 6 are considered in Section 3. Proposition 7 is treated in Section 4. Finally, results of numerical simulations are presented in Section 5.
2 The lower bound on $z^{*}$
As we mentioned in the introduction, the main idea of our proof of Theorem 4 is to construct a feasible solution $x$ to the LP (1) satisfying $\sqrt{2\log(m/n)}\,\langle c,x\rangle\geq 1-\varepsilon$ via an iterative process. The resulting vector $x:=x_{0}+x_{1}+\dots+x_{r}$ will satisfy all the constraints with high probability. To carry this argument over, we start with a number of preliminary estimates concerning subgaussian vectors as well as certain estimates on the condition numbers of random matrices.
We write $[m]$ for the set of integers $\{1,2,\dots,m\}$. Further, given a finite subset of integers $I$, we write $\mathbb{R}^{I}$ for the linear real space of $|I|$–dimensional vectors with components indexed over $I$. Given a set or an event $S$, we will write $\chi_{S}$ for the indicator of the set/event.
2.1 Auxiliary results for subgaussian variables
The following result is known, see for example [50, Section 3.1].
Proposition 9.
Let $V$ be a vector in $\mathbb{R}^{k}$ with mutually independent centered $K$–subgaussian components of unit variance. Then for every $t>0$,
${\mathbb{P}}\big{\{}\big{|}\|V\|_{2}-\sqrt{k}\big{|}\geq t\big{\}}\leq 2\exp(-K_{1}t^{2}),$
where the constant $K_{1}>0$ depends only on $K$.
We recall that, given a standard Gaussian variable $g$,
${\mathbb{P}}\{g\geq\sqrt{t}\}\leq\exp(-t/2),\quad t>0$ (3)
(see, for example, [19, Volume 1, Chapter 7]). The next proposition states that a similar property holds for linear combinations of mutually independent subgaussian variables, as long as the vector of the coefficients is sufficiently incompressible:
Proposition 10 (Asymptotic deviation estimates for linear combinations).
Let $n\to\infty$; for each $n$, let $c=c(n)$ be a non-random unit vector in $\mathbb{R}^{n}$, and let $\delta=\delta(n)\in(0,1)$ be a sequence of numbers such that $\delta n$ converges to infinity. Assume that for every $\nu>0$ and all large $n$, the vector $c(n)$ is $(\nu^{-1}\delta,1-\nu)$–incompressible. Fix any $K\geq 1$, and for each $n$, let $\xi_{i},\xi_{2},\dots,\xi_{n}$ be mutually independent symmetric $K$–subgaussian variables of unit variance. Then for every $\varepsilon>0$ and all large $n$ we have
${\mathbb{P}}\bigg{\{}\sum_{i=1}^{n}c_{i}\xi_{i}\geq(1+\varepsilon)\sqrt{\delta n}\bigg{\}}\leq\exp\big{(}-\delta n/2\big{)}.$
Remark.
Note that, up to the $(1+\varepsilon)$–correction (for arbitrarily small $\varepsilon>0$), the bound in the proposition matches the bound for the standard Gaussians stated above. It is a standard fact in the probability theory that under the assumption $\|c(n)\|_{\infty}=o(1)$, the linear combinations $\sum_{i=1}^{n}c_{i}\xi_{i}$ are asymptotically normally distributed, hence, in view of (3), for every fixed number $s>0$ we have
${\mathbb{P}}\bigg{\{}\sum_{i=1}^{n}c_{i}\xi_{i}\geq\sqrt{s}\bigg{\}}\leq\exp\big{(}-s/2\big{)}+o(1).$
The key input of Proposition 10 is that the “Gaussian–type” estimate also holds true in the regime of larger deviations (in the applications of the proposition, we will choose $\delta n$ roughly of the magnitude of $\log(m/n)$).
Proof of Proposition 10.
Throughout the proof, we assume that $K,\varepsilon>0$ are fixed. Let $\kappa>0$ be a small parameter depending on $K$ and $\varepsilon$, which we will determine below. We start by partitioning the set of indices $[n]$ into two subsets $I=I(n),J=J(n)$ according to magnitude of the absolute values of $c_{i}$’s, namely, we set
$I:=\big{\{}i\leq n:\;|c_{i}|\leq\kappa/\sqrt{\delta n}\big{\}},\quad J:=[n]\setminus I.$
Let $b_{i}$, $i\in I$, be i.i.d symmetric sign variables independent from $\xi_{1},\xi_{2},\dots,\xi_{n}$. Since the random variables $\xi_{i}$’s are symmetrically distributed, the sum $\sum_{i\in I}c_{i}\xi_{i}$ is equidistributed with $\sum_{i\in I}c_{i}b_{i}\xi_{i}$. Note that for each $i\in I$, the variable $\xi_{i}^{2}-1$ is subexponential with the subexponential moment depending only on $K$ (see [50, Chapter 2]). Hence, applying Bernstein’s inequality [50, Section 2.8], we get
${\mathbb{P}}\bigg{\{}\Big{|}\sum_{i\in I}c_{i}^{2}(\xi_{i}^{2}-1)\Big{|}\geq t\bigg{\}}\leq 2\exp\Big{(}-K^{\prime}\min\Big{(}\frac{t^{2}}{\sum_{i\in I}c_{i}^{4}},\frac{t}{\max_{i\in I}c_{i}^{2}}\Big{)}\Big{)},\quad t>0,$
where $K^{\prime}>0$ depends only on $K$. In view of the definition of $I$,
$\sum_{i\in I}c_{i}^{4}\leq\kappa^{2}/(\delta n),\quad\max_{i\in I}c_{i}^{2}\leq\kappa^{2}/(\delta n).$
Thus, assuming that $\kappa=\kappa(K,\varepsilon)$ was chosen sufficiently small and $n$ is large enough,
${\mathbb{P}}\bigg{\{}\sqrt{\sum_{i\in I}c_{i}^{2}\xi_{i}^{2}}\geq 1+\varepsilon/4\bigg{\}}\leq\frac{1}{3}\exp(-\delta n/2)$
(we can in fact provide a much stronger estimate but the one above suffices for our purposes). Further, we use the known fact that for every non-random vector $u\in\mathbb{R}^{I}$,
${\mathbb{P}}\bigg{\{}\sum_{i\in I}u_{i}b_{i}\geq t\bigg{\}}\leq\exp\big{(}-t^{2}/(2\|u\|_{2}^{2})\big{)},\quad t>0$
(see [7]). In particular, whenever $u$ satisfies $\|u|_{2}\leq 1+\varepsilon/4$, and for all large enough $n$, we have
${\mathbb{P}}\bigg{\{}\sum_{i\in I}u_{i}b_{i}\geq(1+\varepsilon/2)\sqrt{\delta n}\bigg{\}}\leq\frac{1}{3}\exp(-\delta n/2).$
Combining the last two estimates, we obtain
$\begin{split}{\mathbb{P}}&\bigg{\{}\sum_{i\in I}c_{i}\xi_{i}\geq(1+\varepsilon/2)\sqrt{\delta n}\bigg{\}}\\ &\leq{\mathbb{P}}\bigg{\{}\sqrt{\sum_{i\in I}c_{i}^{2}\xi_{i}^{2}}\geq 1+\varepsilon/4\bigg{\}}+{\mathbb{P}}\bigg{\{}\sqrt{\sum_{i\in I}c_{i}^{2}\xi_{i}^{2}}\leq 1+\varepsilon/4\quad\mbox{and}\quad\sum_{i\in I}c_{i}b_{i}\xi_{i}\geq(1+\varepsilon/2)\sqrt{\delta n}\bigg{\}}\\ &\leq\frac{2}{3}\exp(-\delta n/2).\end{split}$ (4)
Now, consider the sum $\sum\limits_{i\in J}c_{i}\xi_{i}$. In view of our assumptions on $c$, and the estimate $|J|=O(\delta n)$ which follows from the definition of $J$, the quantity
$\sum\limits_{i\in J}c_{i}^{2}$
converges to zero as $n\to\infty$. Hence, applying Hoeffding’s inequality for subgaussian variables (see [50, Section 2.6]), we obtain for all large enough $n$,
${\mathbb{P}}\bigg{\{}\sum\limits_{i\in J}c_{i}\xi_{i}\geq\frac{\varepsilon}{2}\sqrt{\delta n}\bigg{\}}\leq\frac{1}{3}\exp(-\delta n/2)$ (5)
(again, we can obtain a much stronger estimate but the one above is sufficient). A combination of (4) and (5) completes the proof. ∎
Proposition 9 and a simple union bound argument yields
Corollary 11.
For every $K\geq 1$ there is a number $K_{1}=K_{1}(K)>0$ depending only on $K$ with the following property. Let $k\leq m/e$, let $I\subset[m]$ be a $k$–set, and let $V$ be a random vector in $\mathbb{R}^{[m]\setminus I}$ with mutually independent centered $K$–subgaussian components of unit variance. Then the event
[…]
$\Big{(}\sum_{i\in[m]\setminus I}V_{i}^{2}\,\chi_{\{V_{i}\geq K_{1}\sqrt{\log(m/k)}\}}\Big{)}^{1/2}\leq K_{1}\sqrt{k\log(m/k)}\bigg{\}}$
has probability as least $1-2\big{(}\frac{k}{m}\big{)}^{2k}$.
Proof.
Set
$L:=K_{1}\sqrt{\log(m/k)},$
where $K_{1}=K_{1}(K)\geq 1$ is chosen sufficiently large (the specific requirements on $K_{1}$ can be extracted from the argument below). Since the vector $V$ is subgaussian, every component of $V$ satisfies
${\mathbb{P}}\big{\{}V_{i}\geq L\big{\}}\leq\exp(-K_{2}L^{2})$
for some $K_{2}>0$ depending only on $K$. Hence, the probability that more than $k/2$ components of $V$ are greater than $L$, can be estimated from above by
(${m\choose\lceil k/2\rceil}\exp(-K_{2}L^{2}\,\lceil k/2\rceil)\leq\bigg{(}\frac{em}{\lceil k/2\rceil}\bigg{)}^{\lceil k/2\rceil}\exp(-K_{2}L^{2}\,\lceil k/2\rceil).$ (6)
If $K_{1}$ is chosen so that $\frac{em}{\lceil k/2\rceil}\exp(-K_{2}L^{2})\leq\big{(}\frac{k}{m}\big{)}^{4}$ then the last expression in (6) is bounded above by $\big{(}\frac{k}{m}\big{)}^{2k}$.
Similarly to the above and assuming $K_{1}$ is sufficiently large, in view of Proposition 9, for any choice of a $\lfloor k/2\rfloor$–subset $J\subset[m]\setminus I$, the probability that the vector $(V_{i})_{i\in J}$ has the Euclidean norm greater than $L\sqrt{k}$, is bounded above by
(${m\choose\lfloor k/2\rfloor}^{-1}\Big{(}\frac{k}{m}\Big{)}^{2k}.$
Combining the two observations, we get the result. ∎
Our next observation concerns random matrices with subgaussian entries. Recall that the largest and smallest singular values of an $n\times k$ ($k\leq n$) matrix $M$ are defined by
$s_{\max}(M)=\|M\|:=\sup\limits_{y:\,\|y\|_{2}=1}\|My\|_{2},$
and
$s_{\min}(M):=\inf\limits_{y:\,\|y\|_{2}=1}\|My\|_{2}.$
Proposition 12.
For every $K>0$ there are $K_{1},K_{2}>0$ depending only on $K$ with the following property. Let $\frac{n}{k}\geq K_{1}$ and let $M$ be an $n\times k$ random matrix with mutually independent centered subgaussian entries of unit variance and subgaussian moment at most $K$. Further, let $P$ be a non-random $n\times n$ projection operator of rank $n-1$ i.e $P$ is the orthogonal projection onto an $(n-1)$–dimensional subspace of $\mathbb{R}^{n}$. Then
[…]
Proof.
Let $w$ be the unit vector in the null space of $P$ (note that it is determined uniquely up to changing the sign).
Fix any unit vector $y\in\mathbb{R}^{k}$. Then the random vector $My$ has mutually independent components of unit variance, and is subgaussian, with the subgaussian moment of the components depending only on $K$ (see [50, Section 2.6]). Applying Proposition 9, we get
${\mathbb{P}}\big{\{}\big{|}\|My\|_{2}-\sqrt{n}\big{|}\geq t\big{\}}\leq 2\exp(-K_{1}t^{2}),\quad t>0,$ (7)
where $K_{1}>0$ only depends on $K$. Since the projection operator $P$ acts as a contraction we have $\|PMy\|_{2}\leq\|My\|_{2}$ everywhere on the probability space. On the other hand, $\|PMy\|_{2}\geq\|My\|_{2}-|\langle w,My\rangle|$, where, again according to [50, Section 2.6], $\langle w,My\rangle$ is a subgaussian random variable with the subgaussian moment only depending on $K$, and therefore
${\mathbb{P}}\big{\{}\|My\|_{2}-\|PMy\|_{2}\geq t\big{\}}\leq 2\exp(-K_{2}t^{2}),\quad t>0.$ (8)
Combining (7) and (8), we get
${\mathbb{P}}\big{\{}\big{|}\|PMy\|_{2}-\sqrt{n}\big{|}\geq t\big{\}}\leq 2\exp(-K_{3}t^{2}),\quad t>0,$
for some $K_{3}>0$ depending only on $K$.
The rest of the proof of the proposition is based on the standard covering argument (see, for example, [50, Chapter 4]), and we only sketch the idea. Consider a Euclidean $\delta$–net ${\mathcal{N}}$ on $S^{k-1}$ (for a sufficiently small constant $\delta\in(0,1/2]$) i.e a non-random discrete subset of the Euclidean sphere such that for every vector $u$ in $S^{k-1}$ there is some vector in ${\mathcal{N}}$ at distance at most $\delta$ from $u$. It is known that ${\mathcal{N}}$ can be chosen to have cardinality at most $\big{(}\frac{2}{\delta}\big{)}^{k}$ [50, Section 4.2]. In view of the last probability estimate, the event
$\big{\{}\|PMy\|_{2}-\sqrt{n}\big{|}\leq 0.1\sqrt{n}\mbox{ for all }y\in{\mathcal{N}}\big{\}}$
has probability at least $1-2\big{(}\frac{2}{\delta}\big{)}^{k}\exp(-K_{3}n/100)$. Everywhere on that event, we have (see [50, Section 4.4])
$s_{\max}(PM)=\|PM\|\leq\frac{1}{1-\delta}\max\limits_{y\in{\mathcal{N}}}\|PMy\|_{2}\leq\frac{1.1\sqrt{n}}{1-\delta},$
and
$s_{\min}(PM)\geq\min\limits_{y\in{\mathcal{N}}}\|PMy\|_{2}-\delta\|PM\|\geq 0.9\sqrt{n}-\frac{1.1\delta\sqrt{n}}{1-\delta},$
and, in particular,
$\frac{s_{\max}(PM)}{s_{\min}(PM)}\leq 2\quad\mbox{and}\quad s_{\min}(PM)\geq\sqrt{n}/2,$
assuming that $\delta$ is a sufficiently small positive constant. It remains to note that, as long as the aspect ratio $\frac{n}{k}$ is sufficiently large, the quantity $2\big{(}\frac{2}{\delta}\big{)}^{k}\exp(-K_{3}n/100)$ is exponentially small in $n$. ∎
Remark.
It can be shown that $\frac{s_{\max}(PM)}{s_{\min}(PM)}\leq 1+\varepsilon$ and $s_{\min}(PM)\geq(1-\varepsilon)\sqrt{n}$ with probability exponentially close to one for arbitrary constant $\varepsilon>0$ as long as $K_{1}=K_{1}(\varepsilon)$ is sufficiently large.
Corollary 13.
For every $K>0$ there are $K_{1},K_{2}>0$ depending only on $K$ with the following property. Let $\frac{n}{k}\geq K_{1}$, and let $P$ be a non-random $n\times n$ projection operator of rank $n-1$. Further, let $V_{1},V_{2},\dots,V_{k}$ be independent random vectors in $\mathbb{R}^{n}$ with mutually independent centered $K$–subgaussian components of unit variance. Let $U$ be any non-random vector in $\mathbb{R}^{k}$. Then with probability $1-\exp(-K_{2}n)$ there is the unique vector $y$ in the linear span of $P(V_{1}),P(V_{2}),\dots,P(V_{k})$ such that
$\langle V_{j},y\rangle=U_{j},\quad 1\leq j\leq k,$ (9)
and, moreover, that vector satisfies $\|y\|_{2}\leq 4\|U\|_{2}/\sqrt{n}$.
Proof.
Denote by $M$ the $n\times k$ matrix with columns $V_{1},V_{2},\dots,V_{k}$, and let the random vector $v$ in $\mathbb{R}^{k}$ satisfy $(PM)^{\top}PMv=U$. We then set $y:=PMv$. Note that $y$ is in the linear span of $P(V_{1}),P(V_{2}),\dots,P(V_{k})$ and satisfies (9). Further, we can write
$\|y\|_{2}\leq s_{\max}(PM)\,\|v\|_{2}\leq\frac{s_{\max}(PM)\,\|U\|_{2}}{s_{\min}((PM)^{\top}PM)}=\frac{s_{\max}(PM)\,\|U\|_{2}}{s_{\min}(PM)^{2}}.$
It remains to apply Proposition 12 to get the result. ∎
The next lemma provides a basis for a discretization argument which we will apply in the proof of the theorem.
Lemma 14.
There is a universal constant $K_{1}\geq 1$ with the following property. Let $k\geq 1$. Then there is a set ${\mathcal{N}}_{k}$ of vectors in $\mathbb{R}^{k}$ satisfying all of the following:
  • All vectors in ${\mathcal{N}}_{k}$ have non-negative components taking values in $\frac{1}{\sqrt{k}}\mathbb{N}_{0}$;
  • The Euclidean norm of every vector in ${\mathcal{N}}_{k}$ is in the interval $[1,2]$;
  • For every unit vector $u\in\mathbb{R}^{k}$ there is a vector $y=y(u)\in{\mathcal{N}}_{k}$ such that $u_{j}\leq y_{j}$, $j\leq k$;
  • The size of ${\mathcal{N}}_{k}$ is at most $K_{1}^{k}$.
Proof.
Define ${\mathcal{N}}_{k}$ as the collection of all vectors in $\mathbb{R}^{k}$ with non-negative components taking values in $\frac{1}{\sqrt{k}}\mathbb{N}_{0}$ and with the Euclidean norm in the interval $[1,2]$. Observe that for every unit vector $u\in\mathbb{R}^{k}$, by defining $y$ coordinate-wise as
$y_{j}:=\frac{1}{\sqrt{k}}\,\big{\lceil}\sqrt{k}|u_{j}|\big{\rceil},\quad j\leq k,$
we get a vector with Euclidean norm in the interval $[1,2]$, i.e a vector from ${\mathcal{N}}_{k}$. Thus, ${\mathcal{N}}_{k}$ constructed this way satisfies the first three properties listed above. It remains to verify the upper bound on $|{\mathcal{N}}_{k}|$.
To estimate the size of ${\mathcal{N}}_{k}$ we apply the standard volumetric argument. Observe that $|{\mathcal{N}}_{k}|$ can be bounded from above by the number of integer lattice points in the rescaled Euclidean ball $\sqrt{k}B_{2}^{k}$ of radius $\sqrt{k}$. That, in turn, is bounded above by the maximal number of non-intersecting parallel translates of the unit cube $(0,1)^{k}$ which can be placed inside the ball $2\sqrt{k}B_{2}^{k}$. Comparison of the Lebesgue volumes of $2\sqrt{k}B_{2}^{k}$ and $(0,1)^{k}$ confirms that the latter is of order $\exp(O(k))$, completing the proof. ∎
Equipped with Corollaries 11 and 13, and Lemma 14, we can complete the proof of the theorem.
2.2 Proof of Theorem 4
From now on, we fix a small constant $\varepsilon>0$.
First, suppose that $m=n^{\omega(1)}$. Observe that in this case $\sqrt{2\log(m/n)}=(1-o(1))\sqrt{2\log m}$. Define $\delta=\delta(n):=2(1+\varepsilon/2)(\log m)/n$, so that by the assumptions of the theorem, for every constant $\kappa\in(0,1)$ and all large enough $n$, the cost vectors $c(n)$ are $(\kappa^{-1}\delta,1-\kappa)$–incompressible. Hence, applying Proposition 10, we get for every $i\leq m$ and all large enough $n$,
${\mathbb{P}}\big{\{}(Ac)_{i}\geq(1+\varepsilon/2)\sqrt{\delta n}\big{\}}\leq\exp(-\delta n/2),$
implying
${\mathbb{P}}\big{\{}\|Ac\|_{\infty}\geq(1+\varepsilon/2)\sqrt{\delta n}\big{\}}\leq m\exp(-\delta n/2)=n^{-\omega(1)}.$
We infer that, by considering $x:=\frac{c}{(1+\varepsilon/2)\sqrt{2(1+\varepsilon/2)\log m}}$, the optimal objective value of the LP satisfies
$\sqrt{2\log(m/n)}\,z^{*}\geq\frac{\sqrt{2\log(m/n)}}{(1+\varepsilon/2)\sqrt{2(1+\varepsilon/2)\log m}}>1-\varepsilon$
with probability $1-n^{-\omega(1)}$, and the statement of the theorem follows. In view of the above, to complete the proof it is sufficient to verify the statement under the assumption $m(n)=n^{O(1)}$. In what follows, we assume that $m(n)\leq n^{K^{\prime}}$ for some number $K^{\prime}>0$ independent of $n$.
We shall construct a “good” event of high probability, condition on a realization of the matrix $A$ from that event, and then follow the iterative scheme outlined at the beginning of the section to construct a feasible solution. Set
$x_{0}=x_{0}(n):=\frac{(1-\varepsilon)c}{\sqrt{2\log(m/n)}},$
and define an integer $k_{0}=k_{0}(n)$ as
$k_{0}:=\big{\lceil}\max\big{(}\sqrt{n},m(n/m)^{1+\varepsilon/4}\big{)}\big{\rceil}.$
Definition of a “good” event. Define
[…]
where $\tilde{K}>0$ is a sufficiently large universal constant whose value can be extracted from the argument below. Observe that, by our definition of $k_{0}$ and by our assumption that $\varepsilon$ is small,
$(1+\varepsilon/16)\sqrt{2(1+\varepsilon/16)\log(m/k_{0})}\leq\frac{1-\varepsilon/2}{1-\varepsilon}\sqrt{2\log(m/n)},$
whence for every $i\leq m$, in view of the definition of $x_{0}$,
${\mathbb{P}}\big{\{}(Ax_{0})_{i}\geq 1-\varepsilon/2\big{\}}\leq{\mathbb{P}}\big{\{}(Ac)_{i}\geq(1+\varepsilon/16)\sqrt{2(1+\varepsilon/16)\log(m/k_{0})}\big{\}}.$
Define $\delta=\delta(n):=2(1+\varepsilon/16)\log(m/k_{0})/n$, and observe that, by the assumptions of the theorem, for every constant $\kappa\in(0,1)$ and all large enough $n$, the cost vectors $c(n)$ are $(\kappa^{-1}\delta,1-\kappa)$–incompressible. Hence, we can apply Proposition 10 to get for all large enough $n$,
${\mathbb{P}}\big{\{}(Ac)_{i}\geq(1+\varepsilon/16)\sqrt{2\log(m/k_{0})}\big{\}}\leq\exp\big{(}-\delta n/2\big{)}.$
In particular, the probability that $(Ax_{0})_{i}\geq 1-\varepsilon/2$ for more than $k_{0}$ indices $i$, is bounded above by
(${m\choose k_{0}}\exp\big{(}-\delta nk_{0}/2\big{)}\leq\bigg{(}\frac{em\exp(-\delta n/2)}{k_{0}}\bigg{)}^{k_{0}}=o\bigg{(}\bigg{(}\frac{k_{0}}{m}\bigg{)}^{\Omega(\varepsilon^{2})k_{0}}\bigg{)}.$
On the other hand, for every $t\geq\frac{2\sqrt{k_{0}}}{\sqrt{2\log(m/n)}}$, by Proposition 9 and by the union bound over $k_{0}$–subsets of $[m]$, the probability that the sum of squares of largest $k_{0}$ components of $Ax_{0}$ is greater than $t^{2}$, is bounded above by
(${m\choose k_{0}}\cdot 2\exp(-K_{3}t^{2}\log(m/n))\leq 2\,\Big{(}\frac{em}{k_{0}}\Big{)}^{k_{0}}\exp(-K_{3}t^{2}\log(m/n)),$
where $K_{3}>0$ is a universal constant. Taking $t$ to be a sufficiently large constant multiple of $\sqrt{k_{0}}$ and combining the above estimates, we obtain that the event ${\mathcal{E}}_{c}$ has probability $1-n^{-\omega(1)}$.
Fix for a moment any subset $I\subset[m]$ of size at most $k_{0}$ and at least $\log n$. Further, denote by ${\mathcal{N}}_{I}$ the non-random discrete subset of vectors in $\mathbb{R}^{I}$ with the properties listed in Lemma 14. Denote by $P$ the projection onto the orthogonal complement of the cost vector $c$. Define the event
[…]
such that $\langle y,{\rm row}_{i}(A)\rangle=u_{i}$, $i\in I$; moreover, that vector satisfies:
(a) $\|y\|_{2}\leq 8/\sqrt{n}$;
(b) the number of indices $i\in[m]\setminus I$ with $(Ay)_{i}\geq K^{\prime\prime}\sqrt{\log(m/|I|)}/\sqrt{n}$ is at most $|I|/2$;
$(c)\;\Big{(}\sum_{i\in[m]\setminus I}(Ay)_{i}^{2}\,\chi_{\{(Ay)_{i}\geq K^{\prime\prime}\sqrt{\log(m/|I|)}/\sqrt{n}\}}\Big{)}^{1/2}\leq K^{\prime\prime}\sqrt{|I|\log(m/|I|)}/\sqrt{n}\bigg{\}},$
where $K^{\prime\prime}=K^{\prime\prime}(K)>0$ is a sufficiently large constant depending only on $K$. By combining Corollaries 11 and 13, we obtain that the event ${\mathcal{E}}_{I}$ has probability at least $1-2(|I|/m)^{2|I|}-\exp(-\Omega(n))$, whence the intersection of events $\bigcap\limits_{I\in\mathcal{I}}{\mathcal{E}}_{I}$ taken over the collection $\mathcal{I}$ of all subsets of $[m]$ of size at least $\log n$ and at most $k_{0}$, has probability $1-n^{-\omega(1)}$.
Define the “good” event
${\mathcal{E}}_{good}:={\mathcal{E}}_{c}\;\cap\;\bigcap\limits_{I\in\mathcal{I}}{\mathcal{E}}_{I},$
so that ${\mathbb{P}}({\mathcal{E}}_{good})=1-n^{-\omega(1)}$. Our goal is to show that, conditioned on any realization of $A$ from ${\mathcal{E}}_{good}$, the optimal objective value of the LP is at least $1-\varepsilon$. From now on, we assume that the matrix $A$ is non-random and satisfies the conditions from the definition of ${\mathcal{E}}_{good}$.
Definition of $x_{1}$. Let $I_{0}$ be the $k_{0}$–subset of $[m]$ corresponding to $k_{0}$ largest components of the vector $Ax_{0}$ (we resolve ties arbitrarily). In view of our conditioning on ${\mathcal{E}}_{c}$, for every $i\in[m]\setminus I_{0}$ we have
$\langle{\rm row}_{i}(A),x_{0}\rangle\leq 1-\varepsilon/2.$
Denote by $z_{0}\in\mathbb{R}^{I_{0}}$ the restriction of the vector $Ax_{0}$ to $\mathbb{R}^{I_{0}}$, and let $u_{0}\in{\mathcal{N}}_{I_{0}}$ be a vector which majorizes $z_{0}/\|z_{0}\|_{2}$ coordinate-wise. Note that in view of the definition of ${\mathcal{E}}_{c}$, the Euclidean norm of $z_{0}$ is of order $O(\sqrt{k_{0}})$. Further, let $y_{1}$ be the unique vector in the linear combination of $P({\rm row}_{i}(A))$, $i\in I_{0}$, such that $\langle y_{1},{\rm row}_{i}(A)\rangle=(u_{0})_{i}$, $i\in I_{0}$, satisfying the conditions (a), (b), (c) from the definition of ${\mathcal{E}}_{I_{0}}$. Define $x_{1}:=-\|z_{0}\|_{2}\,y_{1}$, so that $\langle{\rm row}_{i}(A),x_{1}\rangle=-\|z_{0}\|_{2}\,(u_{0})_{i}$, $i\in I_{0}$. Observe that with such a definition of $x_{1}$, the sum $x_{0}+x_{1}$ satisfies
$\langle{\rm row}_{i}(A),x_{0}+x_{1}\rangle=(z_{0})_{i}-\|z_{0}\|_{2}\,(u_{0})_{i}\leq 0,\quad i\in I_{0},$
whereas $\langle x_{0}+x_{1},c\rangle=\langle x_{0},c\rangle=\frac{1-\varepsilon}{\sqrt{2\log(m/n)}}$. Applying the definition of the event ${\mathcal{E}}_{I_{0}}$ and the bound $\|z_{0}\|_{2}=O(\sqrt{k_{0}})$, we get that the vector $x_{1}$ satisfies
The number of indices $i\in[m]\setminus I_{0}$ with $(Ax_{1})_{i}\geq\tilde{K}\sqrt{(k_{0}/n)\log(m/k_{0})}$ is at most $k_{0}/2$;
$\Big{(}\sum_{i\in[m]\setminus I_{0}}(Ax_{1})_{i}^{2}\,\chi_{\{(Ax_{1})_{i}\geq\tilde{K}\sqrt{(k_{0}/n)\log(m/k_{0})}\}}\Big{)}^{1/2}\leq\tilde{K}\sqrt{(k_{0}^{2}/n)\log(m/k_{0})},$
where $\tilde{K}=\tilde{(}K)>0$ depends only on $K$. We will simplify the conditions on $x_{1}$. First, observe that in view of the definition of $k_{0}$ and since $m=n^{O(1)}$, we can assume that $\tilde{K}\sqrt{(k_{0}/n)\log(m/k_{0})}\leq\varepsilon/16$. Further, we can estimate $\tilde{K}\sqrt{(k_{0}^{2}/n)\log(m/k_{0})}$ from above by $\sqrt{\lfloor k_{0}/2\rfloor}$. Thus, the vector $x_{1}$ satisfies
The number of indices $i\in[m]\setminus I_{0}$ with $(Ax_{1})_{i}\geq\varepsilon/16$ is at most $\lfloor k_{0}/2\rfloor$;
$\Big{(}\sum_{i\in[m]\setminus I_{0}}(Ax_{1})_{i}^{2}\,\chi_{\{(Ax_{1})_{i}\geq\varepsilon/16\}}\Big{)}^{1/2}\leq\sqrt{\lfloor k_{0}/2\rfloor}.$
Definition of $x_{2},x_{3},\dots$ The vectors $x_{2},x_{3},\dots$ are constructed via an inductive process. Let $s:=\lceil\log_{2}k_{0}-\frac{1}{4}\log_{2}n\rceil$. We define numbers $k_{1},k_{2},\dots,k_{s}$ recursively by the formula
$k_{j}:=\lfloor k_{j-1}/2\rfloor,\quad 1\leq j\leq s.$
With $n$ large, we then have
$k_{s}=n^{1/4\pm o(1)}.$
Further, assuming the vector $x_{j}$ has been defined, we let $I_{j}$ be a $k_{j}$–subset of $[m]\setminus I_{j-1}$ corresponding to $k_{j}$ largest components of $Ax_{j}$ (with ties resolved arbitrarily).
Let $11\leq h\leq s-$ 1, and assume that a vector $x_{h}$ has been constructed that satisfies
The number of indices $i\in[m]\setminus I_{h-1}$ with $(Ax_{h})_{i}\geq(3/4)^{h-1}\varepsilon/16$ is at most $k_{h}$;
$\Big{(}\sum_{i\in[m]\setminus I_{h-1}}(Ax_{h})_{i}^{2}\,\chi_{\{(Ax_{h})_{i}\geq(3/4)^{h-1}\varepsilon/16\}}\Big{)}^{1/2}\leq\sqrt{k_{h}}$
(observe that $x_{1}$ satisfies the above assumptions, which provides a basis of induction). We denote by $z_{h}\in\mathbb{R}^{I_{h}}$ the restriction of the vector $Ax_{h}$ to $\mathbb{R}^{I_{h}}$ (note that $z_{h}$ has the Euclidean norm at most $\sqrt{k_{h}}$ by the induction hypothesis), and let $u_{h}\in{\mathcal{N}}_{I_{h}}$ be a vector which majorizes $z_{h}/\|z_{h}\|_{2}$ coordinate-wise. Similarly to the above, we let $y_{h+1}$ be the unique vector in the linear span of $P({\rm row}_{i}(A))$, $i\in I_{h}$, such that $\langle{\rm row}_{i}(A),y_{h+1}\rangle=(u_{h})_{i}$, $i\in I_{h}$, and satisfying the conditions (a), (b), (c) from the definition of ${\mathcal{E}}_{I_{h}}$, and define $x_{h+1}:=-\|z_{h}\|_{2}\,y_{h+1}$, so that $\langle{\rm row}_{i}(A),x_{h+1}\rangle=-\|z_{h}\|_{2}\,(u_{h})_{i}$, $i\in I_{h}$. By the definition of ${\mathcal{E}}_{I_{h}}$, the vector $Ax_{h+1}$ satisfies
The number of indices i[m]Ih with (Axh+1)iK~khlog(m/kh)/n is at most $\begin{split}&\mbox{The number of indices $i\in[m]\setminus I_{h}$ with $(Ax_{h+1})_{i}\geq\tilde{K}\sqrt{k_{h}\log(m/k_{h})}/\sqrt{n}$ is at most $k_{h+1}$;}\\ &\Big{(}\sum_{i\in[m]\setminus I_{h}}(Ax_{h+1})_{i}^{2}\,\chi_{\{(Ax_{h+1})_{i}\geq K^{\prime\prime}\sqrt{k_{h}\log(m/k_{h})}/\sqrt{n}\}}\Big{)}^{1/2}\leq\tilde{K}\sqrt{k_{h}^{2}\log(m/k_{h})}/\sqrt{n},\end{split}$ (10)
for some constant $\tilde{K}=\tilde{K}(K)>0$. In is not difficult to check that, with our definition of $k_{1},k_{2},\dots$ and the assumption $m=n^{O(1)}$ and as long as $n$ is large, the quantity $\tilde{K}\sqrt{k_{h}\log(m/k_{h})}/\sqrt{n}$ is (much) less than $(3/4)^{h}\varepsilon/16$; further, $\tilde{K}\sqrt{k_{h}^{2}\log(m/k_{h})}/\sqrt{n}$ is dominated by $\sqrt{k_{h+1}}$. Thus, $x_{h+1}$ satisfies
The number of indices $i\in[m]\setminus I_{h}$ with $(Ax_{h+1})_{i}\geq(3/4)^{h}\varepsilon/16$ is at most $k_{h+1}$;
$\Big{(}\sum_{i\in[m]\setminus I_{h}}(Ax_{h+1})_{i}^{2}\,\chi_{\{(Ax_{h+1})_{i}\geq(3/4)^{h}\varepsilon/16\}}\Big{)}^{1/2}\leq\sqrt{k_{h+1}},$
completing the induction step.
Restoring feasibility. The above process produces vectors $x_{1},x_{2},\dots,x_{r}$, where $x_{r}$ satisfies, in view of (10) and the choice of $r$,
$\Big{(}\sum_{i\in[m]\setminus I_{r-1}}(Ax_{r})_{i}^{2}\,\chi_{\{(Ax_{r})_{i}\geq(3/4)^{r-1}\varepsilon/16\}}\Big{)}^{1/2}<\varepsilon/16,$
so that, in particular, $(x_{r})_{i}\leq\varepsilon/16$ for every $i\in[m]\setminus I_{r-1}$. Note also that for every $h\leq r-1$ and $i\in I_{h}$, we have
$\langle{\rm row}_{i}(A),x_{h}+x_{h+1}\rangle\leq 0.$
To complete the proof, we have to verify that the sum $x:=x_{0}+x_{1}+\dots+x_{r}$ is a feasible solution to the linear program. Fix any index $i\in[m]$. Let $\chi_{0}(i)$ be the Boolean indicator of the expression “$\langle{\rm row}_{i}(A),x_{0}\rangle\geq 1-\varepsilon/2$”, for each $11\leq h\leq r-$ 1, let $\chi_{h}(i)$ be the indicator of “$\langle{\rm row}_{i}(A),x_{h}\rangle\geq(3/4)^{h-1}\varepsilon/16$”, and set $\chi_{r}(i):=0$. Note that, in view of the definition of the sets $I_{0},I_{1},\dots$, $\chi_{h}(i)=1$ only if $\chi_{h+1}(i)=0$. The set $\{0,1,\dots,r\}$ can be partitioned into consecutive subsets $S_{1}<S_{2}<\dots<S_{p}$ in such a way that
  • The size of each set $S_{\ell}$ is either $1$ or $2$, and
  • Whenever $\chi_{h}(i)=1$ for some $h$, the index $h$ belongs to a $2$–subset from the partition, and it is the smaller of the two indices in that subset.
We have
$\langle{\rm row}_{i}(A),x\rangle=\sum_{r=1}^{p}\sum_{h\in S_{r}}\langle{\rm row}_{i}(A),x_{h}\rangle.$
Fix for a moment any $r\leq p$ and consider the corresponding sum $\sum_{h\in S_{r}}\langle{\rm row}_{i}(A),x_{h}\rangle$. First, assume that $S_{r}$ is a singleton, and let $i$ be its element. By our construction of the partition, we have $\chi_{h}(i)=0$.
  • If $i=0$ then
    $\langle{\rm row}_{i}(A),x_{i}\rangle\leq 1-\varepsilon/2;$
  • If $11\leq i\leq r-$ 1 then
    $\langle{\rm row}_{i}(A),x_{i}\rangle\leq(3/4)^{h-1}\varepsilon/16;$
  • If $i=r$ then
    $\langle{\rm row}_{i}(A),x_{i}\rangle\leq\varepsilon/16.$
Next, assume that the size of $S_{r}$ is two, and represent the subset as $S_{r}=\{h,h+1\}$. By the construction of the vectors $x_{h}$ and $x_{h+1}$, we then have
$\langle{\rm row}_{i}(A),x_{h}+x_{h+1}\rangle\leq 0.$
Combining the two possible cases and summing over $r$, we get
$\langle{\rm row}_{i}(A),x\rangle\leq 1-\varepsilon/2+\sum_{h=1}^{\infty}(3/4)^{h-1}\varepsilon/16+\varepsilon/16\leq 1,$
and the proof is finished.
3 Applications of Theorem 4
Proof of Corollary 5.
Assume that the cost vectors $c=c(n)$ satisfy $\sqrt{\log(m/n)}\|c\|_{\infty}=0$. Fix any constant $\kappa\in(0,1)$, and let $I=I(n)$ be any subset of $[n]$ of size at most $\kappa^{-1}\log(m/n)$. In view of the assumptions on $c$, we have
$\sum_{i\in I}c_{i}^{2}\leq\|c\|_{\infty}^{2}|I|\leq\kappa^{-1}\log(m/n)\|c\|_{\infty}^{2}\longrightarrow 0.$
Since the choice of the subsets $I(n)$ was arbitrary, the last relation implies that $c$ is $(\kappa^{-1}\log(m/n),1-\kappa)$–incompressible for all large $n$. It remains to apply Theorem 4. ∎
Proof of Corollary 6.
In this proof, we assume that for each $n$, $c=c(n)$ is a uniform random unit vector in $\mathbb{R}^{n}$, and that $c(n)$, $n\in\mathbb{N}$, are mutually independent and independent from $\{A(n),\;n\in\mathbb{N}\}$. Recall that the mean width of the polyhedron $P=P(n)=\{x\in\mathbb{R}^{n}:\;Ax\leq{\bf 1}\}$ is defined as
${\mathcal{W}}(P)=2\,{\mathbb{E}}_{c}\,\max\{\langle c,x\rangle:\;x\in P\},$
where ${\mathbb{E}}_{c}$ denotes the expectation taken with respect to $c=c(n)$. It will be most convenient for us to prove the statement by contradiction. Namely, assume that there is a small positive number $\delta>0$ such that
${\mathbb{P}}\big{\{}\liminf\limits_{n\to\infty}\sqrt{2\log(m/n)}{\mathcal{W}}(P(n))\leq 2-\delta\big{\}}\geq\delta.$
Denote by $\mathcal{F}$ the $\sigma$–field generated by the random matrices $\{A(n),\;n\in\mathbb{N}\}$. Observe that the event
:=${\mathcal{E}}:=\big{\{}\liminf\limits_{n\to\infty}\sqrt{2\log(m/n)}{\mathcal{W}}(P(n))\leq 2-\delta\big{\}}$
is $\mathcal{F}$–measurable. Condition for a moment on any realization of the polyhedra $P(n)$, $n\in\mathbb{N}$, from ${\mathcal{E}}$, and let $(n_{k})_{k=1}^{\infty}$ be an increasing sequence of integers such that $\sqrt{2\log(m/n_{k})}{\mathcal{W}}(P(n_{k}))\leq 2-\delta/2$ for every $k$. The condition ${\mathcal{W}}(P(n_{k}))\leq\frac{2-\delta/2}{\sqrt{2\log(m/n_{k})}}$ implies
$\frac{1-\delta/4}{\sqrt{2\log(m/n_{k})}}$ $\geq{\mathbb{E}}_{c}\,\max\{\langle c(n_{k}),x\rangle:\;x\in P(n_{k})\}$
$\geq\frac{1-\delta/8}{\sqrt{2\log(m/n_{k})}}\,{\mathbb{P}}_{c}\bigg{\{}\max\{\langle c(n_{k}),x\rangle:\;x\in P(n_{k})\}\geq\frac{1-\delta/8}{\sqrt{2\log(m/n_{k})}}\bigg{\}},$
and hence
${\mathbb{P}}_{c}\bigg{\{}\max\{\langle c(n_{k}),x\rangle:\;x\in P(n_{k})\}<\frac{1-\delta/8}{\sqrt{2\log(m/n_{k})}}\bigg{\}}\geq 1-\frac{1-\delta/4}{1-\delta/8},$
where ${\mathbb{P}}_{c}$ is the conditional probability given the realization of $P(n)$, $n\in\mathbb{N}$, from ${\mathcal{E}}$. By the Borel–Cantelli lemma, the last assertion yields
${\mathbb{P}}_{c}\bigg{\{}\liminf\limits_{n\to\infty}\sqrt{2\log(m/n)}\max\{\langle c(n),x\rangle:\;x\in P(n)\}\leq 1-\delta/8\bigg{\}}=1.$
Removing the conditioning on ${\mathcal{E}}$, we arrive at the estimate
${\mathbb{P}}\bigg{\{}\liminf\limits_{n\to\infty}\sqrt{2\log(m/n)}\max\{\langle c(n),x\rangle:\;x\in P(n)\}\leq 1-\delta/8\bigg{\}}\geq\delta.$ (11)
In the second part of the proof, we will show that the assertion (11) leads to contradiction. Fix for a moment any $\kappa\in(0,1)$. We claim that $c(n)$ is $(\kappa^{-1}\log(m/n),1-\kappa)$–incompressible for every large enough $n$ with probability $1-n^{-\omega(1)}$. Indeed, the standard concentration inequality on the sphere (see, in particular, [50, Chapter 5]) implies that for every choice of a non-random $\lfloor\kappa^{-1}\log(m/n)\rfloor$–subset $I\subset[n]$,
${\mathbb{P}}\bigg{\{}\sqrt{\sum_{i\in I}c_{i}^{2}}\geq\sqrt{\frac{K_{1}|I|}{n}}+t\bigg{\}}\leq 2\exp(-K_{2}nt^{2}),\quad t>0,$
for some universal constants $K_{1},K_{2}>0$. Taking the union bound, we then get
${\mathbb{P}}$$\bigg{\{}\sqrt{\sum_{i\in I}c_{i}^{2}}\geq\sqrt{\frac{K_{1}|I|}{n}}+t\;\;\mbox{for some $\lfloor\kappa^{-1}\log(m/n)\rfloor$--subset $I$}\bigg{\}}$
$\leq 2{n\choose\lfloor\kappa^{-1}\log(m/n)\rfloor}\exp(-K_{2}nt^{2}),\quad t>0,$
Our assumption $m=\exp(o(n))$ implies that
$\sqrt{\frac{K_{1}\,\lfloor\kappa^{-1}\log(m/n)\rfloor}{n}}\leq\kappa/2\quad\mbox{for all large enough $n$.}$
For all such $n$, the last estimate with $t:=\kappa/2$ yields
${\mathbb{P}}\bigg{\{}\sqrt{\sum_{i\in I}c_{i}^{2}}\geq\kappa\;\;\mbox{for a $\lfloor\kappa^{-1}\log(m/n)\rfloor$--subset $I$}\bigg{\}}\leq 2{n\choose\lfloor\kappa^{-1}\log(m/n)\rfloor}\exp(-K_{2}n\kappa^{2}/4).$ (12)
It remains to observe that as long as $m=\exp(o(n))$, we have
(${n\choose\lfloor\kappa^{-1}\log(m/n)\rfloor}=\exp(o(n)),$
to that the expression on the right hand side of (12) is bounded above by $2\exp(-\Omega(n))$. This verifies the claim.
With the claim above confirmed, it remains to apply Theorem 4. From the theorem, we obtain that
$\liminf\limits_{n\to\infty}\sqrt{2\log(m/n)}\max\{\langle c,x\rangle:\;x\in P(n)\}\geq 1$
with probability one. However, that contradicts (11), completing the proof. ∎
4 The outer radius of random Gaussian polytopes
In this section, we prove the bound on the outer radius of random polytopes stated in the introduction. We would like to emphasize that, unlike Theorem 4, the argument of this section is entirely based on standard considerations. We provide the proof for completeness.
The following approximation of the standard Gaussian distribution is well known (see, for example, [19, Volume I, Chapter VII, Lemma 2]):
Lemma 15.
Let $g$ be a standard Gaussian variable. Then for every $t>0$,
$\Big{(}\frac{1}{t}-\frac{1}{t^{3}}\Big{)}\exp(-t^{2}/2)\leq{\mathbb{P}}\{g\geq t\}\leq\frac{1}{t}\exp(-t^{2}/2).$
Let $G$ be a standard Gaussian random vector in $\mathbb{R}^{m}$, and $G_{(1)}\leq G_{(2)}\leq\dots\leq G_{(m)}$ be the non-decreasing rearrangement of its components. Let $21\leq k\leq m/$ 2. Then, from the above lemma, for every $t>0$ we have
${\mathbb{P}}\big{\{}G_{(m-k)}\leq t\big{\}}\leq{m\choose k}{\mathbb{P}}\{g\leq t\}^{m-k}\leq\Big{(}\frac{em}{k}\Big{)}^{k}\Big{(}1-\Big{(}\frac{1}{t}-\frac{1}{t^{3}}\Big{)}\exp(-t^{2}/2)\Big{)}^{m-k}.$ (13)
As a corollary, we obtain the following deviation estimate:
Corollary 16.
Fix any $\varepsilon\in(0,1)$. Let sequences of positive integers $m(n)$, $n\geq 1$, and $k(n)$, $n\geq 1$, satisfy $m=\omega(n)$, $k=\omega(1)$, and $k=o(m)$. Further, for each $n$, let $G=G(n)$ be a standard Gaussian vector in $\mathbb{R}^{m}$. Then
${\mathbb{P}}\big{\{}G_{(m-k)}\leq(1-\varepsilon)\sqrt{2\log(m/k)}\big{\}}=\exp\Big{(}-\Big{(}\frac{m}{k}\Big{)}^{\Omega(1)}k\Big{)}.$
Proof.
Applying (13), we get
${\mathbb{P}}$$\big{\{}G_{(m-k)}\leq(1-\varepsilon)\sqrt{2\log(m/k)}\big{\}}$
$\leq\Big{(}\frac{em}{k}\Big{)}^{k}\bigg{(}1-\Big{(}\frac{1}{(1-\varepsilon)(2\log(m/k))^{1/2}}-\frac{1}{(1-\varepsilon)^{3}(2\log(m/k))^{3/2}}\Big{)}\Big{(}\frac{k}{m}\Big{)}^{(1-\varepsilon)^{2}}\bigg{)}^{m-k}$
$=\exp\Big{(}-\Big{(}\frac{m}{k}\Big{)}^{\Omega(1)}k\Big{)},$
implying the claim. ∎
Proof of Proposition 7.
Fix any $\varepsilon>0$. Observe that the statement is equivalent to showing that with probability $1-n^{-\omega(1)}$ the matrix $A$ has the property that for every $y\in S^{n-1}$ there is $i\leq m$ with $(Ay)_{i}\geq(1-\varepsilon)\sqrt{2\log(m/n)}$.
Let $k$ be the largest integer bounded above by
$m\Big{(}\frac{n}{m}\Big{)}^{1-\varepsilon/8}.$
Define $\delta:=\exp(-(m/n)^{K})$, where $K>0$ is a sufficiently small constant depending only on $\varepsilon$ (the value of $K$ can be extracted from the argument below). Let ${\mathcal{N}}$ be a Euclidean $\delta$–net in $S^{n-1}$ i.e a subset of the sphere such that for every $y\in S^{n-1}$ there is $y^{\prime}\in{\mathcal{N}}$ satisfying $\|y-y^{\prime}\|_{2}\leq\delta$. It is known that ${\mathcal{N}}$ can be chosen to have cardinality at most $\big{(}\frac{2}{\delta}\big{)}^{n}$ (see [50, Chapter 4]).
Fix for a moment any $y^{\prime}\in{\mathcal{N}}$, and observe that, in view of Corollary 16 and the definition of $k$, we have
$(Ay^{\prime})_{i}\geq(1-\varepsilon/2)\sqrt{2\log(m/n)}\quad\mbox{for at least$k$ indices $i\leq m$}$ (14)
with probability at least $1-\exp\big{(}-\big{(}\frac{m}{k}\big{)}^{\Omega(1)}k\big{)}\geq 1-\exp(-k)$. Therefore, assuming that $K$ is sufficiently small and taking the union bound over all elements of ${\mathcal{N}}$, we obtain that with probability $1-n^{-\omega(1)}$, (14) holds simultaneously for all $y^{\prime}\in{\mathcal{N}}$.
Condition on any realization of $A$ such that the last event holds and, moreover, such that $\|A\|\leq 2\sqrt{m}$ (recall that the latter occurs with probability $1-\exp(-\Omega(m))$; see [50, Chapter 4]). Assume for a moment that there is $y\in S^{n-1}$ such that $(Ay)_{i}\leq(1-\varepsilon)\sqrt{2\log(m/n)}$ for all $i\leq m$. In view of the above, it implies that there is a vector $y^{\prime}\in{\mathcal{N}}$ at distance at most $\delta$ from $y$ such that
$(A(y^{\prime}-y))_{i}\geq\frac{1}{2}\varepsilon\sqrt{2\log(m/n)}$
for at least $k$ indices $i\leq m$. Hence, the spectral norm of $A$ can be estimated as
$\|A\|\geq\frac{\varepsilon\sqrt{2k\log(m/n)}}{2\delta}>2\sqrt{m},$
contradicting the bound $\|A\|\leq 2\sqrt{m}$. The contradiction shows that for every $y\in S^{n-1}$ there is $i\leq m$ such that $(Ay)_{i}\geq(1-\varepsilon)\sqrt{2\log(m/n)}$. The result follows. ∎
5 Numerical experiments
Our results described in the previous sections naturally give rise to the questions: (a) How close to each other are the asymptotic estimates of the optimal objective value and empirical observations?, (b) What is the asymptotic distribution and standard deviation of $z^{*}$?, and (c) How does the algorithm in the proof of Theorem 4 perform in practice? We discuss these questions in the following subsections, and make conjectures while providing numerical evidence. We use Gurobi 10.0.1 for solving instances of the LP (1), and set all Gurobi parameters to default.
5.1 Magnitude of the optimal objective value
Note that although the asymptotic lower bound of Theorem 4 is the best possible distribution-independent estimate of $z^{*}$, the result does not rule out the scenario when for some subgaussian distributions of the matrix entries, the optimal objective value either has no limit or has a limit which is strictly greater than in the Gaussian setting. It is not difficult to see that the estimate of Proposition 7 fails for random matrices $A(n)$ with i.i.d Rademacher variables; in fact, in the Rademacher case the outer radius of the polyhedron is at least $1$ everywhere on the probability space which can be confirmed immediately by considering standard basis vectors as feasible solutions. At the same time, we expect that for sufficiently incompressible cost vectors the lower bound on $z^{*}$ in Theorem 4 predicts the correct asymptotics.
Below, we investigate the quality of approximation of the optimal objective value $z^{*}$ by the asymptotic bound $l_{b}:=(2\log(m/n))^{-\frac{1}{2}}$ given in Theorem 4. The empirical outcomes are obtained through multiple runs of the LP (1) under various choices of parameters. We consider two distributions of the entries of $A$: either $A$ is standard Gaussian or its entries are i.i.d Rademacher (±$\pm 1$) random variables. In either case and for different choices of $m$, we sample the random LP (1) taking the sample size $50$ and letting $c$ be the vector of i.i.d Rademacher variables rescaled to have unit Euclidean norm. The cost vector is generated once and remains the same for each of the $50$ instances of the LP within a sample.
[…]
1000 50 0.40853 0.50626 23.92
2000 50 0.36816 0.44119 19.83
6000 50 0.32317 0.37256 15.28
10000 50 0.30719 0.35176 14.50
20000 50 0.28888 0.32473 12.41
Table 1: Asymptotic versus empirically observed optimal objective value in the Gaussian setting (the sample size = $50$)
As is shown in Tables 1 and 2, for a fixed value of $n=50$, as the number of constraints $m$ progressively increases in magnitude, the relative gap between $l_{b}$ and the sample mean $\hat{\mu}$ of the optimal objective values, defined as
$\left(\frac{|l_{b}-\hat{\mu}|}{l_{b}}\times 100\right)\%,$
decreases.
[…]
1000 50 0.40853 0.50369 23.29
2000 50 0.36816 0.43801 18.97
6000 50 0.32317 0.37200 15.11
10000 50 0.30719 0.35220 14.65
20000 50 0.28888 0.32856 13.73
Table 2: Asymptotic versus empirically observed optimal objective value in the Rademacher setting (the sample size = $50$)
The empirical results enable us to formulate the following conjecture:
Conjecture 17 (Upper bound on $z^{*}$ for uniform random cost vectors).
Let $m$, $n$, $A(n)$ be as in Theorem 4, and for each $n$ let $c=c(n)$ be uniform random unit cost vector independent from $A(n)$. Then $\limsup\limits_{n\to\infty}\sqrt{2\log(m/n)}z^{*}\leq 1$ almost surely.
A counterpart of the above conjecture deals with the mean width of the feasible regions:
Conjecture 18 (Upper bound on the mean width).
Let $m$, $n$, $A(n)$ be as in Theorem 4, and for each $n$ let $P=P(n):=\{x\in\mathbb{R}^{n}:\;Ax\leq{\bf 1}\}$. Then $\limsup\limits_{n\to\infty}\sqrt{2\log(m/n)}\,{\mathcal{W}}(P)\leq 2$ almost surely.
5.2 Structural assumptions on the cost vector
In this subsection, we carry out a numerical study of the technical conditions on the cost vectors from Theorem 4.
[…]
1 0.429907 15.08
2 0.454357 10.25
3 0.459387 9.25
4 0.479007 5.38
5 0.481839 4.82
6 0.486141 3.97
7 0.479635 5.25
8 0.493168 2.58
9 0.490829 3.04
10 0.498035 1.62
Table 3: The sample mean of the objective value corresponding to a given value of the parameter $k$. The random matrix $A$ has dimensions $m=1000$ and $n=50$, with each entry equidistributed with the product of Bernoulli($1/2$) and N(00,$2$) variables. The sample size = $50$. The third column of the table is the relative gap between the obtained sample mean and the value of the sample mean from Table 1 for the same $m$, $n$.
Roughly speaking, those conditions stipulate that the cost vector has significantly more than $\log(m/n)$ “essentially non-zero” components. Since Theorem 4 is an asymptotic result, it is not at all clear what practical assumptions on the cost vectors should be imposed to guarantee that the resulting optimal objective value is close to the asymptotic lower bound.
In the following experiment, we consider a random coefficient matrix $A$ with i.i.d subgaussian components equidistributed with the product of Bernoulli($1/2$) and N(00,$2$) random variables. Note that with this definition the entries have zero mean and unit variance. Further, we consider a family of cost vectors $c(n,k)$ parametrized by an integer $k$, so that the vector $c(n,k)$ has $k$ non zero components equal to $1/\sqrt{k}$ each, and the rest of the components are zeros. For a given choice of $k$, we sample the LP (1) with the above distribution of the entries and the cost vector $c(n,k)$. Our goal is to compare the resulting sample mean of the optimal objective value to the sample mean obtained in the previous subsection for the same matrix dimensions and for the strongly incompressible cost vector. Our empirical results show that, as long as $k$ is small (so that $c(n,k)$ is very sparse), the value of the sample mean differs significantly from the one in Table 1. On the other hand, for $k$ sufficiently large the relative gap between the sample means becomes negligible. The results are presented in Table 3 below.
5.3 Limiting distribution of the optimal objective value
Recall that, given mutually independent standard Gaussian variables, their maximum converges (up to appropriate rescaling and translation) to the Gumbel distribution as the number of variables tends to infinity. If the vertices of the random polyhedron $P=\{x\in\mathbb{R}^{n}:\;Ax\leq{\bf 1}\}$ were behaving like independent Gaussian vectors (with a proper rescaling) then the limiting distribution of $z^{*}$ for any given fixed cost vector would be asymptotically Gumbel. Our computations suggest that in fact the limiting distribution is Gaussian.
((a))
((b))
Figure 1: The frequency histogram and the empirical cumulative distribution function of the optimal objective values for the Gaussian matrix $A$ with $m=1000$, $n=50$ (the sample size = $1000$). The KS test results: KS statistic = $0.0232$, $p$-value = $0.6453$.
((a))
((b))
Figure 2: The frequency histogram and the empirical cumulative distribution function of the optimal objective values for the Rademacher matrix $A$ with $m=1000$, $n=50$ (the sample size = $1000$). The KS test results: KS statistic = $0.0219$, $p$-value = $0.7161$.
In our numerical experiments, we consider Gaussian and Rademacher random coefficient matrices of dimensions $m=1000$ and $n=50$. The sample size in either case is taken to be $1000$. As in the previous experiment, we take $c$ to be a random vector with i.i.d Rademacher components rescaled to have the Euclidean norm one. We generate the cost vector once so that it is the same for every LP from a sample. We employ the Kolmogorov–Smirnov (KS) test to compare the sample distribution of $z^{*}$ to Gaussian of a same mean and variance. In either case, the obtained p-value exceeds the conventional significance level of 0.05, indicating lack of substantial evidence to refute the null hypothesis that the data is normally distributed. Further, visual representation of the frequency histogram and the empirical cumulative distribution function of the optimal values shown in Figures 1 and 2 validate this conjecture.
The next conjecture summarizes the experimental findings:
Conjecture 19 (Limiting distribution of $z^{*}$).
Let $m$, $n$, $A(n)$ be as in Theorem 4, and for each $n$ let $c=c(n)$ be uniform random unit cost vector independent from $A(n)$. Then the sequence of normalized random variables
$\frac{z^{*}-{\mathbb{E}}\,z^{*}}{\sqrt{{\rm Var}\,z^{*}}}$
converges in distribution to the standard Gaussian.
5.4 Standard deviation of the optimal objective value
As in the previous two numerical experiments, in this one we consider two types of the entries distributions: Gaussian and Rademacher. The cost vector is generated according to the same procedure as above. We sample the LP (1) taking the sample size to be $50$, and compute the sample standard deviation $\hat{\sigma}$ (the square root of the sample variance) of the optimal objective value. Based on our numerical studies we speculate that standard deviation of $z^{*}$ is roughly of the order $1/\sqrt{m}$, at least in the setting where $m$ is very large compared to $n$. Indeed, upon examination of the last column in Tables 4 and 5, we observe a consistent phenomenon: the quantities are of order $1$ for various choices of $m$ and $n$.
$m$ $n$ $l_{b}$ $\hat{\sigma}$ $\hat{\sigma}\times\sqrt{m}$
1000 50 0.408539 0.02420 0.765
2000 50 0.368161 0.01728 0.773
4000 50 0.337791 0.01333 0.843
6000 50 0.323170 0.01142 0.885
8000 50 0.313877 0.01059 0.947
10000 50 0.307196 0.01027 1.027
1000 100 0.465991 0.03014 0.953
2000 100 0.408539 0.01592 0.712
4000 100 0.368161 0.01172 0.742
6000 100 0.349456 0.00990 0.767
1000 150 0.528257 0.03192 1.010
2000 150 0.441515 0.015066 0.674
4000 150 0.391745 0.01155 0.731
6000 150 0.368161 0.011850 0.918
Table 4: Sample standard deviation of the optimal objective value with Gaussian matrix $A$ (the sample size = $50$)
$m$ $n$ $l_{b}$ $\hat{\sigma}$ $\hat{\sigma}\times\sqrt{m}$
1000 50 0.408539 0.020420 0.645
2000 50 0.368161 0.017393 0.777
4000 50 0.337791 0.011962 0.756
6000 50 0.323170 0.010207 0.790
8000 50 0.313877 0.010333 0.924
10000 50 0.307196 0.008997 0.899
1000 100 0.465991 0.023097 0.730
2000 100 0.408539 0.014648 0.655
4000 100 0.368161 0.010436 0.660
6000 100 0.349456 0.011791 0.913
1000 150 0.528257 0.032600 1.030
2000 150 0.441515 0.015466 0.691
4000 150 0.391745 0.011030 0.697
6000 150 0.368161 0.009153 0.708
Table 5: The sample standard deviation of the optimal objective value with Rademacher matrix $A$ (the sample size = $50$)
5.5 An algorithm for finding a near-optimal feasible solution
As previously mentioned, the proof of Theorem 4 involves an iterative projection process designed to restore feasibility from an initial point. This process involves discretizing a continuous space spanned by violated constraints at each iteration (a covering argument) as a means to apply probabilistic union bound argument. A direct implementation of this process is rather heavy, however, a closely related and at the same time computationally efficient alternative would be an instance of the classical Kaczmarz block-iterative projection method. Specifically, we focus on algorithms presented in [18] or their randomized counterpart in [38] and [53]. In our context, the blocks correspond to the rows of the matrix $A$ that are violated by the initial point and its updates at each iteration. We did not attempt to further optimize the algorithm in any other aspects including the number of iterations it takes to produce a feasible solution. Still, we think it is of interest to consider a practical realization of the algorithm and test it on instances of LP (1). Table 6 shows the results of running the algorithm for various parameters $m$ and $n$. For each instance of the LP, $r$ is the the number of iterations required to restore feasibility, $x_{0}=(2\log(m/n))^{-\frac{1}{2}}c$ is the initial point and $x$ is the output of the Algorithm 1. Note that the obtained perturbations to $x_{0}$ are in the orthogonal complement of the cost vector $c$ and as a result $\|x_{0}\|_{2}=z(x)$. The numbers $|I_{0}|$ and $|I_{1}|$ in the last two columns of the Table are defined in the Algorithm 1, specifically, $|I_{0}|$ and $|I_{1}|$ are the number of violated constraints by $x_{0}$ at iteration 1 and $x_{0}+x_{1}$ at iteration 2, respectively. Note that these numbers are rather small compared to the size of the LP instance. We note that in our experiment, the algorithm did not converge for the $20000\times 50$ instance of the LP, which we attribute to the random nature of the problem. At the same time, the Algorithm converged fast due to small number of violations for all other given instances.
Input: $A_{m\times n}$: standard Gaussian random matrix,
             $c$: random cost vector uniformly distributed on the sphere
Output: $x$: near-optimal feasible solution to LP (1)
1 $x\leftarrow(2\log(m/n))^{-\frac{1}{2}}c$;
2 $j\leftarrow 0$;
3 $\epsilon\leftarrow 0.1$;
4 while $x$ is not feasible do
5       $j\leftarrow j+1$;
6       Find the set of constraints $I_{j-1}$ that are violated or nearly violated by $x$, i.e all indices $i$ with $\langle{\rm row}_{i}(A),x\rangle>1-\epsilon$;
7       Compute vector $b$ where $b_{i}=1-\epsilon-\langle{\rm row}_{i}(A),x\rangle$ for all $i\in I_{j-1}$;
8       Project the constraints into the space $c^{\perp}$  (i.e define $v_{i}={\rm row}_{i}(A)-\langle{\rm row}_{i}(A),c\rangle c$ for all $i\in I_{j-1}$) ;
9       Find $x_{j}$ as a linear combination of $v_{i}$’s, $i\in I_{j-1}$, so that the constraints violated by $x$ are repaired (specifically, solve the linear system $M^{T}Mu=b$ for $u$, then solve $x_{j}=Mu$ where $M$ is the matrix with columns $v_{i}$);
10       $\epsilon\leftarrow\epsilon*0.1$;
11       $x\leftarrow x+x_{j}$;
12
13 end while
return $x$
Algorithm 1 Algorithm for finding near-optimal feasible solution to LP (1), an instance of block-iterative Kaczmarz projection method
$m$ $n$ $r$ $\|x_{0}\|_{2}=z(x)$ $|I_{0}|$ $|I_{1}|$
1000 50 2 0.408539 11 1
1000 100 2 0.465991 28 4
2000 50 2 0.368161 6 1
2000 100 2 0.408539 23 1
4000 50 2 0.337791 13 2
4000 100 2 0.368161 30 1
6000 50 2 0.323170 22 21
6000 100 2 0.349456 31 8
8000 50 2 0.313877 22 6
8000 100 2 0.337791 21 5
10000 50 2 0.307196 19 1
10000 100 1 0.329505 34 0
20000 100 1 0.307196 29 0
40000 50 1 0.273493 12 0
40000 100 2 0.288881 34 3
60000 50 2 0.265558 16 2
60000 100 1 0.279576 36 0
80000 50 1 0.260329 21 0
80000 100 2 0.273493 36 1
100000 5 2 0.256479 28 5
100000 100 2 0.269040 35 1
Table 6: Numerical results for the Algorithm 1 for various $m$ and $n$ ($r$ = the number of iterations to restore the feasibility).
6 Conclusion
In this paper, we study the optimal objective value $z^{*}$ of a random linear program where the entries of the $m\times n$ coefficient matrix $A$ have subgaussian distributions, the variables $x_{1},\dots,x_{n}$ are free and the right hand side of each constraint is $1$. We provide sharp asymptotic lower bound $(2\log(m/n))^{-\frac{1}{2}}$ for $z^{*}$ in the regime $n\to\infty$, $\frac{m}{n}\to\infty$. This lower bound provides quantitative information about the geometry of the random polyhedron defined by the feasible region of the random LP (1), specifically about its spherical mean width. The computational experiments support our theoretical guarantees and give insights into several open questions regarding the asymptotic behaviour of the random LP. The connection between the proof of the main result and convex feasibility problems allows us to provide a fast algorithm for obtaining a near optimal solution in our randomized setting. Possible future directions of research are provided as conjectures in the computational experiments section. The random LP (1) is a step towards studying models with inhomogeneous coefficient matrices and establishing stronger connections to practical applications.